please write with hand draw

Q: Suppose a 3 year bond with a 6% coupon rate that was purchased for $760 and had a promised yield of 8%. Suppose that interest rates increased and the price of the bond declined. Displeased, you sold the bond for 798.8 after having owned it for 1 year. What should be the realized yield ?
Answer #1

Face value of Bond = $1,000

Annual coupon rate = 6%

Coupon amount(C) = 6%*1000 = $60

Purchase price(P) = $760

Sell Price(S) = $798.8

Holding period(n) = 1

Realized yield (r) would be:

d Ta 1-, (+5) = u/t +S

798.8 +60 1/1 r = 760 ) - 1

T= 1.13-1

r = 13.0%